OpenStudy (anonymous):

1. Many people know that the weight of an object varies on different planets, but did you know that the weight of an object on Earth also varies according to the elevation of the object? In particular, the weight of an object follows this equation: , where C is a constant, and r is the distance that the object is from the center of Earth. a. Solve the equation for r. b. Suppose that an object is 150 pounds when it is at sea level. Find the value of C that makes the equation true. (Sea level is 3,963 miles from the center of the Earth.) c. Use the value of C you found in the previous que

OpenStudy (saifoo.khan):

Online college? right.

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

I am no good at math lol

OpenStudy (saifoo.khan):

Cool, many people doing online college come for help here! U will surely get help here! :)

OpenStudy (anonymous):

that is nice cause I don't understand it and need to badly I have a 60% in my math class right now

OpenStudy (anonymous):

where is the equation?

OpenStudy (anonymous):

the weight of an object follows this equation:......... ,

OpenStudy (saifoo.khan):

LOL, help from here can make u get 80+.. @kumar, u have to make the equation!

OpenStudy (saifoo.khan):

I've done this same problem like 1000 times! :(

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

oh no wow that is alot

OpenStudy (saifoo.khan):

LOL.

OpenStudy (saifoo.khan):

kumar, u solving?

OpenStudy (saifoo.khan):

??

OpenStudy (anonymous):

I think he left

OpenStudy (saifoo.khan):

yea, he did.

OpenStudy (saifoo.khan):

let's do it step by step..

OpenStudy (anonymous):

k cause they want me to show work and I wanna learn how if that's ok

OpenStudy (anonymous):

khan i dont think so, there must be an equation. read it carefully.

OpenStudy (saifoo.khan):

Kumar, i have done this like 1000 time, we have to make the equation!

OpenStudy (anonymous):

\[W=Cr ^{-2}\]

OpenStudy (anonymous):

this is it I am sure

OpenStudy (saifoo.khan):

In particular, the weight of an object follows this equation: w= Cr^-2. where C is a constant, and r is the distance that the object is from the center of the earth. a. Solve the equation w=Cr^-2 for r. Multiply both sides by r^2 to get: wr^2 = C r^2 = C/w r = sqrt[C/w]

OpenStudy (anonymous):

It says I can use 1 mile=5,280 as feet for the conversions

OpenStudy (saifoo.khan):

Right

OpenStudy (saifoo.khan):

i got the doc file with the answers, which they give that to you!

OpenStudy (anonymous):

oh i c

OpenStudy (saifoo.khan):

let me find that, just a sec.

OpenStudy (anonymous):

k

OpenStudy (saifoo.khan):

u want the explaination or the answers right away?

OpenStudy (anonymous):

which is easier

OpenStudy (anonymous):

I have alot more homework to do question wise I do want to learn though

OpenStudy (saifoo.khan):

lol, i can do that.. but Openstudy people take it the bad way!

OpenStudy (anonymous):

I really need to pass this class and I am scared I am not lol I have bombed almost everything so I finally thought I would try this

OpenStudy (saifoo.khan):

copy this asap.

OpenStudy (anonymous):

I just did

OpenStudy (saifoo.khan):

;]

OpenStudy (anonymous):

ty

OpenStudy (saifoo.khan):

uw.

OpenStudy (saifoo.khan):

please hit good answer.

OpenStudy (anonymous):

Done!

OpenStudy (saifoo.khan):

are u in Univ of Pheonix?

OpenStudy (saifoo.khan):

Hi Bonez.

OpenStudy (anonymous):

The next one can I put here

OpenStudy (saifoo.khan):

put that on left.

OpenStudy (anonymous):

yep